HRMS, 28th Colloquium, Dijon, France, 28 Aug - 1 Sep, 2023

A NEW TREATMENT OF THE 2ν_{3} (A_{1}) INFRARED BAND OF METHYL BROMIDE CH_{3}Br AROUND 1213 cm^{−1}

\underline{HAMID NAJIB} , Laboratoire des Energies Renouvelables et Environnement, Département de Physique, Faculté des Sciences, Université Ibn Tofaïl, BP: 133,

Kénitra, MOROCCO

The high-resolution Fourier transform infrared spectrum of methyl bromide CH_{3}Br has been reinvestigated in the v_{3} = 2 vibrational excited state around 1213 cm^{−1}, for both^{79}Br and^{81}Br isotopologues. Thanks to new accurate K-dependent and J-dependent ground state parameters, up to sextic centrifugal distortion constants, 420 new infrared transitions of the 2ν_{3} band have been assigned for CH_{3}^{79}Br and CH_{3}^{81}Br, extending the rotational quantum number values up to K_{max} = 15 and J_{max} = 60. For each isotopic species, a standard deviation of 2 x 10^{−4} cm^{−1} has been obtained for about 700 lines of 2ν_{3}. The new parameter values of the v_{3} = 2 excited state are in good agreement with ones determined previously by infrared spectroscopy, but significantly more accurate.

1Hamid Najib J. Mol. Spectrosc., 370, (2020)111297.

p-number: p070  Submitted on Sun, 11 Jun 2023 17:10:13 +0200